OpenStudy (anonymous):

Evaluate the limit: Limit of the function (7/x - (7 - |x|)) as x approaches 0 from the left side (0-)

OpenStudy (anonymous):

If you are looking at it from the left, then x will be less than 0. Then it will be negative, so if you put in that, you get: (7/x - (-7/x) = (7/x + 7/x) = 14/x taking the limit of function (14/x) as x approaches 0 from the left side DNE right?

OpenStudy (anonymous):

But then when I enter this answer "DNE" into my electronic assignment, it says that this answer is wrong.

OpenStudy (anonymous):

\[\frac{7}{x} - \left( 7 - |x|\right)\]?

OpenStudy (anonymous):

No it is: (7/x) - (7/|x|)